OpenStudy (anonymous):

if cos inverse x/2+cos inverse y/3=z ,then prove that 9x square-12xy cos z +4y square=36 sin square z

OpenStudy (unklerhaukus):

\[\arccos\left(\frac x2\right)+\arccos \left(\frac y3\right)=z\]

OpenStudy (unklerhaukus):

\[9x ^2-12xy \cos \left(z +4y ^2\right)=36 \sin ^2( z)\]

OpenStudy (unklerhaukus):

like this?

OpenStudy (anonymous):

yes . it is..

OpenStudy (unklerhaukus):

i bet you have to take a square root at some stage

OpenStudy (anonymous):

do arc and inverse,both signify the same thing?

OpenStudy (unklerhaukus):

yer

OpenStudy (anonymous):

okay. can you help me in solving out it.at staring of steps?

OpenStudy (unklerhaukus):

i dont know what to do

OpenStudy (anonymous):

okay.thanks